MAIN MENU Risk and Rates of Return 1 My Home < Back to Assignment 9 Courses Attempts: Average: /1 CENGAGE UNLIMITED 6. Proble8.14
Answer #1

Current Portfolio Beta = 1.02

Weight of each stock in portfolio = 1/20 = 0.05

Beta of sell stock = 1.0

Beta of New stock = 1.48

Thus,

New Portfolio Beta would be:

= 1.02 -0.05 + 1.0 +0.05*1.48

= 1.044

=1.04

Hope it will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.